OpenStudy (anonymous):

Solve the linear system: x+2y=2 x-2y=6

OpenStudy (anonymous):

x+2y = 2 ---- call this equation (1) x-2y = 6 ----call this equation (2) use (1) + (2) you get 2x = 8 so x = 8/4 = ? after you get the value of x, substitute x into either one of the two equations, you will have the value of y

OpenStudy (anonymous):

2x=8 so x = 8/2 = ? is what @linkinage meant

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

I guess the 8/2 = ?part is the "guiding the user to the answer" :P lol

OpenStudy (wolf1728):

You have to solve for each unknown by eliminating them one at a time. Add equation A and B A) x+2y=2 B) x-2y=6 2x = 8 so x =4 Try solving for y.

OpenStudy (anonymous):

do we subtract now to solve for y?

OpenStudy (anonymous):

there are several ways to do this. by subtracting two equations you can get the value of y as well. XD

OpenStudy (wolf1728):

Now that x has been eliminated, solving for y is easier. Substitute 'x' into either equation A or B will easily solve for 'y' So let's pick equation A and substitute the 'x' with 4 x+2y=2 4 + 2y = 2 Now can you solve for 'y'?

OpenStudy (anonymous):

6y=2 y=3?

OpenStudy (wolf1728):

No 4 + 2y = 2 Subtract 4 from both sides

OpenStudy (anonymous):

y=-2?

OpenStudy (wolf1728):

I'm guessing you did this 4 -4 + 2y = 2 -4 2y = -2 divide both sides by 2

OpenStudy (anonymous):

y=-1

OpenStudy (wolf1728):

You've got it !!

OpenStudy (anonymous):

thank you!!

OpenStudy (wolf1728):

u r welcome :-)